OpenStudy (anonymous):

how can you convert -0.03 into a fraction? Without a calculator.

OpenStudy (anonymous):

-3/100

OpenStudy (anonymous):

sorry the 3 is repeating

OpenStudy (anonymous):



OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

1/3 = 0.33333... 1/30 = 0.03333...

OpenStudy (anonymous):

You move the decimal to the left. When you divide the number by 10

OpenStudy (anonymous):

\[-\overline {.3}=-\frac{1}{3}\]

OpenStudy (anonymous):

-1/30

OpenStudy (anonymous):

I got negative -1/30 but dont know how to get to that without a calculator.. i did the other two using a formula, but I dont what I am doing wrong.

OpenStudy (anonymous):

here is what you can do. forget the minus sign, it is a distraction. we will put it in at the end

OpenStudy (anonymous):

-0.3333... = -1/3 so divide both side by 10 -0.03333... = -1/30

OpenStudy (anonymous):

put \[x=.03333...\] \[10x=.33333\] \[9x=.3333...-.033333...=0.3\] \[x=\frac{.3}{9}=\frac{3}{90}=\frac{1}{30}\]

OpenStudy (anonymous):

actually moneybird method is slicker, but it requires knowing one fact to start, namely that \[\frac{1}{3}=\overline{.3}\]if you have another problem where you do not know such a thing, you have to use a different method

OpenStudy (anonymous):

for example if you want \[\overline{.12}\] you can write \[x=\overline{.12}\] \[100x=12\overline{.12}\] \[99x=12\] \[x=\frac{12}{99}=\frac{4}{33}\]

OpenStudy (anonymous):

later when this becomes a bit more sophisticated you can sum a geometric series, but for now this gimmick should always work

OpenStudy (anonymous):

yeah that is the formula i used for the other two and got the correct answers and I tried doing the .03 repeating but i am not getting - 1/30

OpenStudy (anonymous):

How do you make the fraction symbol?

OpenStudy (anonymous):

\{a}{b}

OpenStudy (anonymous):

if you want to know any code you see, you can right click on it to see it

OpenStudy (anonymous):

\sum_{k=1}^{\infty}\frac{1}{2^k} \[\sum_{k=1}^{\infty}\frac{1}{2^k}\]

OpenStudy (anonymous):

frac{frac{1}{2}}{1-frac{1}{2}}

OpenStudy (anonymous):

\[\frac{\frac{1}{2}}{1-\frac{1}{2}}\]

OpenStudy (anonymous):

= 1

OpenStudy (anonymous):

there you go. now have fun. there are lots of nice tricks. main reason i started here was to learn latex. now i am hooked

OpenStudy (anonymous):

it seems kind of hard. Before I just press the equation button.

OpenStudy (anonymous):

thanks for tryin to help me